What is the domain and range of the exponential function f(x) = 2^x


1
Expert's answer
2021-01-29T05:08:51-0500

The domain of "f(x)=2^x" is "(-\infty,+\infty)" .

The range of "f(x)=2^x" is "(0,+\infty)". The range doesn't include 0.